Please help to show the workings.

Blossom Corp issued 10-year bonds four years ago with a coupon rate of 9.560 percent. At the time of issue, the bonds sold at par. Today bonds of similar risk and maturity must pay an annual coupon of 8.54 percent to sell at par value. Assuming semiannual coupon payments.

(a)What is the bond’s yield to maturity? (Round answer to 2 decimal places, e.g. 15.25.)

(b) What is the current market price of the firm’s bonds? (Round answer to 2 decimal places, e.g. 15.25.)
Answer #1

a)Since given here after 4 years the bond to sell at par then the coupon rate is 8.54% and this is the YTM of bond

b)please use below formulae to find the price of bond

Price=(coupon*(1-((1+i)^-n))/i)+(issue price*(1+i)^-n)
Coupon=maturity value*coupon rate
=1000*(9.56%/2)=47.80
i=8.54%/2%
n=(10-4)*2=12 period
issue price=1000
substituting in formuale we get it as 1113.55